OpenStudy (anonymous):

5y+30 5 complete the factoring

OpenStudy (anonymous):

is that 5y + 30 or + 305

OpenStudy (anonymous):

whats with the arbitrary 5. whats going on there.

OpenStudy (anonymous):

5y+30 would be 5(y+6)

OpenStudy (anonymous):

5y+30 =5( ) complete the factoring

myininaya (myininaya):

\[5y+30=5 \cdot y+5 \cdot 6 =5(y+6)\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

please answer